OpenStudy (anonymous):

Find z*w. z=5(cos35+isin35) and w=2(cos40+isin40).

OpenStudy (loser66):

we have a formula to find it out. It says z*w = number in the front of z* number in the front of w *(cos (angle of z + angle of w) + i sin (angle of z + angle of w)) show my your understanding

OpenStudy (anonymous):

Would it be 10(cos75+isin75)? (75 being in degrees)

OpenStudy (loser66):

Proud of you!! :)

OpenStudy (anonymous):

Thanks, could you help me with writing in a+bi form?

OpenStudy (loser66):

we have formula to do it, too z= r , theta so, from yours r = 10, theta = 75 the formula to write the complex form is r = x +iy where x = r cos (theta) y = r sin (theta) show me your work

OpenStudy (loser66):

I take the result as z , it is not original z.

OpenStudy (anonymous):

The expression I need to write in a+bi form is (sqrt3 + i)^5.

OpenStudy (loser66):

ohhh, new problem?

OpenStudy (anonymous):

Yes, sorry

OpenStudy (loser66):

let say the number is z = (sqrt3 +i)^5

OpenStudy (loser66):

is it not just open the bracket? you have the form (a + bi )already? where a = sqrt 3, and b =1??

OpenStudy (loser66):

but I don't think we have to do it. it forms the form you need

OpenStudy (anonymous):

That's why I am confused. The 4 answers to choose from are 16-16i(sqrt3) -16(sqrt3)+16i 16(sqrt3)+16i -16(sqrt3)-16i

OpenStudy (loser66):

one way to do, just (sqrt 3 +i)(sqrt3+i)(sqrt3+i)(sqrt3+i)(sqrt3+i) and open the first 2 terms = 3 +2sqrt3*i +i^2 = 2+ 2sqrt3*i open the next 2 terms, get the same so far, we have (2+2sqrt3*i)(2+2sqrt3 *i)(sqrt3 +i) next, the first 2 terms of the new one is..... if you know how to use Pascal triangle , the problem will be easier to solve but if you don't , just count your fingers as what I did, hehehe